WARNING: This web is not used anymore. Please use PDBService.MinuteS10Aug06 instead!

Minutes 10 Aug 06

  • Phone:
    • TRIUMF: Denise
    • FNAL: Eric
    • RAL: Nuala, Gordon
    • BNL Iris
    • GridKA: Doris, Silke, Andreas
    • ATLAS: Sasha

  • CERN:
    • 3D: Dirk, Eva, Luis, Zbigniew
    • COOL: Andrea
    • CMS: Zhen
    • ATLAS: Florbella, Gancho
    • ARDA: Birger

  • General announcements:
    • Reorganization of the wiki page
      • links to the monitoring tools (OEM, new web tool prototype)
      • technical documentation
      • please send the correspondence between the databases installed and the experiments to Eva and Dirk
      • feedback on the wiki pages is welcome (send it to Dirk)
      • experiment tests wiki pages available, please take over
    • Eric will send the URL for the CMS squid tests to Dirk
    • LHCC review: Gordon has been invited to give a presentation
    • Priority: Streams Setup for the experiments

* Round table - site status:

    • ASCG: cluster reinstallation ongoing, no date foreseen
    • BNL (Iris): working with Gancho on the Streams setup for ATLAS. OEM problem, suggested to send an email to Dirk Geppert
    • CNAF: cluster up and available in OEM, second cluster to be ready, Dirk will contact Barbara
    • IN2P3: dbas on holidays till 22th Aug
    • GridKA (Doris): Streams setup ready, throughput tests ongoing. OEM setup: agent installed
    • TRIUMF (Denise): hw has been requested
    • PIC: Dirk will meet Delfino

  • Experiment tests:
    • ATLAS: there is a problem with BNL disk space requirement
    • LHCb: sites have been contacted
    • Integration databases are being prepared: archivelog mode + patches related to Streams
    • Frontier - CMS test: status report during the next meeting

  • Backup and Recovery Tests:
    • Scenarios and Test description available in wiki (Gordon and Eva)
    • Proposed to start with the population scripts used during the throughput tests
    • Eva will setup the environment for the tests
    • Destination sites: RAL (Gordon) and CNAF (no confirmed)

  • LFC tests to be discussed with Barbara and LFC people (Dirk)

  • Next meeting: Thrusday 17, focus: Frontier
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-08-14 - EvaDafonte
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback